### v2

# FlowDreamer: Exploring High Fidelity Text-to-3D Generation Via Rectified Flow

Abstract
Recent advances in text-to-3D generation have made significant progress. In particular, with the pretrained diffusion models, existing methods predominantly use Score Distillation Sampling (SDS) to train 3D models such as Neural Radiance Fields (NeRF) and 3D Gaussian Splatting (3D GS). However, a hurdle is that they often encounter difficulties with over-smoothing textures and over-saturating colors. The rectified flow model – which utilizes a simple ordinary differential equation (ODE) to represent a linear trajectory – shows promise as an alternative prior to text-to-3D generation. It learns a time-independent vector field, thereby reducing the ambiguity in 3D model update gradients that are calculated using time-dependent scores in the SDS framework. In light of this, we first develop a mathematical analysis to seamlessly integrate SDS with rectified flow model, paving the way for our initial framework known as Vector Field Distillation Sampling (VFDS). However, empirical findings indicate that VFDS still results in over-smoothing outcomes. Therefore, we analyze the grounding reasons for such a failure from the perspective of ODE trajectories. On top, we propose a novel framework, named FlowDreamer, which yields high-fidelity results with richer textual details and faster convergence.The key insight is to leverage the coupling and reversible properties of the rectified flow model to search for the corresponding noise, rather than using randomly sampled noise as in VFDS. Accordingly, we introduce a novel Unique Couple Matching (UCM) loss, which guides the 3D model to optimize along the same trajectory. Our FlowDreamer is superior in its flexibility to be applied to both NeRF and 3D GS. Extensive experiments demonstrate the high-fidelity outcomes and accelerated convergence of FlowDreamer. Moreover, we highlight the intriguing open questions, such as initialization challenges in NeRF and sampling techniques, to benefit the research community. For more details, please visit our project page at https://vlislab22.github.io/FlowDreamer.
Figure 1: FlowDreamer uses a pretrained rectified flow model to generate high-fidelity results from text prompts. It can generate not only highly realistic objects, such as guns and shoes, but also fantastical ones, such as dragon heads.
1 Introduction
3D generation enjoys broad applications in diverse fields, such as the Metaverse, games, education, architecture design, and films, and has attracted significant research interest recently (Xie et al., 2024; Wang et al., 2024b; Tang et al., 2024; Poole et al., 2022; Chen et al., 2023a; Lin et al., 2023; Jain et al., 2022; Tang et al., 2023; Jiang et al., 2024). Text-to-3D generation – which generates 3D contents from user-input text – has emerged as one of the promising 3D generation paradigms due to its ease of use (Wang et al., 2023; Yi et al., 2023; Wang et al., 2024a; Nichol et al., 2022; Jun & Nichol, 2023).
Recently, with the advances of text-to-2D image synthesis techniques based on the diffusion models, text-to-3D generation also undergoes a surge of research interest. A seminal work, DreamFusion (Poole et al., 2022) sets a cornerstone by proposing Score Distillation Sampling (SDS) to address this issue by leveraging pretrained text-to-image diffusion model, to train Neural Radiance Fields (NeRF)Mildenhall et al. (2021). It has been rapidly evolved to 3D Gaussian Splatting (3D GS) (Kerbl et al., 2023; Tang et al., 2023; Yi et al., 2023) for faster training and rendering.
Despite the success, some works, e.g., Lin et al. (2023); Poole et al. (2022); Zhu et al. (2023) unveil that SDS suffers from issues, such as over-smoothing textures and over-saturating colors. For this reason, some attempts, e.g.,Wang et al. (2024a); Liang et al. (2023); Wu et al. (2024) improve SDS from different perspectives. For example, Prolificdreamer (Wang et al., 2024a) introduces variational score distillation (VSD), which models 3D parameters as random variables to distill 3D assets, however, it requires significantly more time to optimize. Consistent3D (Wu et al., 2024) designs a consistency distillation sampling method to train the 3D model. Nevertheless, the quality improvements are still limited. Luciddreamer (Liang et al., 2023) proposes interval score matching (ISM) loss in the diffusing trajectory, but the loss drops many terms with the same scale.
Recently, flow matching approaches (Liu et al., 2022; Lipman et al., 2022) pave new ways for fast and high-quality generation. Among them, rectified flow model (Liu et al., 2022; Esser et al., 2024; Liu et al., 2023) uses a simple ordinary differential equation (ODE) to represent a straight trajectory. It trains a time-independent vector field, indicating that the direction of the vector field remains constant. Whereas diffusion trajectories result in curved trajectories(Lipman et al., 2022), score(Song et al., 2020b) is time-dependent, meaning that SDS optimizing over uniformly sampled values of $t$ can produce different gradient directions. Owning to its merits, we interestingly find that it could serve as an alternative prior for text-to-3D generation.
Figure 2: An example of over-smoothing results.
In light of this, we first develop a mathematical analysis to seamlessly integrate SDS with rectified flow model. This enables us to build up an initial framework, named as Vector Field Distillation Sampling (VFDS). However, empirical results demonstrate that VFDS still leads to over-smoothing textures, see Figure 2. To this end, we further analyze the grounding reasons for such a failure from the perspective of ODE trajectories. This way, we find that, as VFDS randomly samples noise, it leads to multiple ODE trajectories in nearly the same images, i.e., from camera poses with mild differences, see Figure 4. This causes inconsistent update directions, leading to over-smoothing textural issues.
Buttressed by the analysis, we propose a novel framework, named FlowDreamer, which yields high-fidelity results with rich textual details and faster convergence. The key idea of FlowDreamer is to leverage the coupling and reversible properties of rectified flow model. Importantly, the reversible property is explored to search the corresponding noise from the image while the coupling property ensures that the corresponding noise is unique. For formality, we define it as a push-backward process to avoid the aforementioned over-smoothing issue caused by multiple ODE trajectories and better make our update directions consistent. Empirical experiments show that the push-backward process is efficient as three Euler discretization steps are sufficient for it to achieve plausible performance, see Figure 1.Accordingly, we propose a novel Unique Couple Matching (UCM) loss that guides the 3D model to learn the same trajectory. Our FlowDreamer also enjoys high flexibility as it can be applied to either NeRF or 3D GS generation settings.
We conduct extensive experiments under diverse generation settings, demonstrating high-fidelity results with rich details and faster convergence of FlowDreamer, as shown in Figure 1. When exploring the application of FlowDreamer to NeRF, we observe some interesting phenomena. Moreover, we identify some intriguing open questions. First is the initialization problem that emerges when applying FlowDreamer to NeRF. This is because the distribution of the initialized image from NeRF is undefined in the Rectified flow diffusion space. We use a warm-up strategy to mitigate this issue. Secondly, the push-backward process with different numbers of function evaluations (NFE) and sampling methods can generate some interesting results.
In summary, our major contributions are as follows:
  • We are the first to explore a new direction by leveraging the rectified flow model as an alternative prior for text-to-3D generation. We accordingly build a mathematical analysis to adapt SDS to rectified flow model, paving the way for an initial framework – VFDS. Empirical results demonstrate that VFDS still leads to over-smoothing. We further analyze the underlying reasons for this issue from the perspective of ODE trajectories.
  • Based on the analysis, we further propose a text-to-3D framework, FlowDreamer, with a novel UCM loss. The loss is build opon the push-backward process to search for corresponding noise, rather than using randomly sampled noise in VFDS.
  • Extensive experiments in both NeRF and 3D GS generation settings demonstrate high-fidelity results with rich details and faster convergence for our FlowDreamer. We also identify some interesting open questions, such as initialization issues for NeRF and sampling techniques in the noise search process.
2 Related Works
Text-to-3D generation. It aims to create 3D assets from user-input text. Dreamfusion (Poole et al., 2022) proposes Score Distillation Sampling (SDS) that leverages the pretrained diffusion models to train a NeRF. However, SDS suffers from issues such as over-smoothing textures, low resolution, slow convergence, multi-faced problem(Wang et al., 2024a; Lin et al., 2023; Poole et al., 2022), etc. Magic3D (Lin et al., 2023) designs a coarse-to-fine two-stage training pipeline and changes the 3D model to DMtet (Shen et al., 2021) to improve the resolution of the generated 3D results. Later on, some works (Tang et al., 2023; Yi et al., 2023; Liang et al., 2023; Chen et al., 2023b; Jiang et al., 2024; Li et al., 2024; Jiang & Wang, 2024) take 3D GS as the 3D model for faster training and rendering. Recently, to overcome the multi-face problem, some works (Shi et al., 2023; Wang & Shi, 2023; Tang et al., 2024) fine-tune the pretrained diffusion models to generate multi-view images.
Design variant of SDS loss. To overcome the issue of over-smoothing textural issues, some attempts (Wang et al., 2023; 2024a; Liang et al., 2023; Wu et al., 2024; Zhu et al., 2023; Katzir et al., 2023; Yu et al., 2023) focus on designing different SDS losses. For example, Wang et al. (2023) proposes Score Jacobian Chaining, which applies the chain rule to the estimated score to enhance generation quality. Prolificdreamer (Wang et al., 2024a) proposes VSD to model 3D parameters as random variables to distill 3D assets. Although with improved quality, it requires a much longer time to optimize. Consistent3D (Wu et al., 2024) designs a consistency distillation sampling method to train the 3D model. Nevertheless, the quality improvements are limited. Luciddreamer (Liang et al., 2023) proposes ISM loss in the diffusing trajectory, but the loss drops many terms with the same scale. These methods build loss towards either DDPM(Ho et al., 2020) or DDIM (Song et al., 2020a) models. By contrast, we propose a novel UCM loss that is build opon the push-backward process to search for corresponding noise, rather than using randomly sampled noise with rectified flow-based models. Our UCM demonstrates high-fidelity results with rich details and faster convergence under either NeRF or GS generation settings.
Diffusion and Flow based models. Recent advances in text-to-image generation have witnessed a significant progress. Diffusion models define a process, which progressively converts a distribution of training data to pure Gaussian noise. By learning the reverse process, one can sample data following the distribution. DDPM (Ho et al., 2020) employs a Markov chain to achieve the above process, while DDIM (Song et al., 2020a) proposes to reduce the iteration step by maintaining the marginal distribution hence free from the restriction of Markov property. The diffusion process can be essentially modeled as stochastic differential equations (SDE) (Song et al., 2020b) or ordinary stochastic equations (ODE) (Lipman et al., 2022). On the other hand, flow-base models (Liu et al., 2022; Lipman et al., 2022; Liu et al., 2023) pave new ways for faster and higher-quality generation.
Recently, rectified flow (Liu et al., 2022) – one of the ODE methods – defines a simple process, optimizing the trajectories in diffusion space to be as straight as possible. We are the first to explore a new direction by leveraging the rectified flow model as an alternative prior for text-to-3D generation. We propose a novel UCM loss, built upon the push-backward process to search for corresponding noise, Extensive experiments in both NeRF and 3D GS generation settings demonstrate higher-fidelity results with richer details and faster convergence.
3 FlowDreamer
3.1 VFDS: SDS in the Lens of Rectified flow
Adapting SDS to the rectified flow framework. We first briefly introduce the rectified flow (Liu et al., 2022). Let $\pi_{1}$ and $\pi_{0}$ denote Gaussian distribution $\mathcal{N}(\mathbf{0},\mathbf{I})$ and data distribution, respectively. $\epsilon$ and $x_{0}$ are respectively sampled from $\pi_{1}$ and $\pi_{0}$. Rectified flow defines the forward process as:
$x_{t}=t\epsilon+(1-t)x_{0},t\in[0,1]$ (1)
Figure 3: (a): An illustration of the reversible and coupling properties of the rectified flow model. The coupling property indicates that $\epsilon$ is connected to $x$ by a single trajectory and does not cross other trajectories. The rectified flow trajectory is bidirectional, indicating that it is reversible. (b) An illustration of the UCM loss. Given a rendered image, the push-backward process is used to search for #$\#_{\phi}[x]$, followed by interpolation and then gradient calculation.
Accordingly, the reverse process follows the Ordinary Differential Equation (ODE) to map $\epsilon$ to $x_{0}$.
$dx_{t}=v_{\phi}(x_{t},t)dt,t\in[0,1],$ (2)
where the velocity $v_{\phi}$ is estimated by a learnable network $\phi$. The model is trained as follows:
$\mathcal{L}_{\mathrm{rflow}}(\phi,\mathbf{x})=\mathbb{E}_{x_{0}\sim p_{0},\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I}),t\sim U[0,1]}\left[w(t)\|(\epsilon-x_{0})-v_{\phi}(x_{t},t)\|_{2}^{2}\right],$ (3)
where $w(t)$ is a time-dependent weighting function, $U[0,1]$ denotes the uniform distribution within $[0,1]$. Because the rectified flow model is an ODE model, it has reversible and coupling properties. The $\epsilon$ from the Gaussian noise distribution is uniquely coupled with the $x$ from the data distribution. Moreover, the rectified flow trajectory is bidirectional, indicating that it is reversible, as shown in Figure 3(a). Specifically,
1) Reversible property: The $\epsilon$ from the Gaussian noise distribution can map to $x$, while $x$ from the data distribution can also map to $\epsilon$.
2) Coupling property: The $\epsilon$ is determined, and the $x$ generated by the same rectified flow model is unique. Conversely, the generated $\epsilon$ is unique for a given $x$.
Now, we elucidate how to adapt SDS to the rectified flow to build an initial framework, called Vector Field Distillation Sampling (VFDS). The loss, denoted as $\mathcal{L}_{\mathrm{VFDS}}$, can be written as:
$\nabla_{\theta}\mathcal{L}_{\mathrm{VFDS}}(\phi,x=g(\mathbf{\theta,c}))=\mathbb{E}_{\epsilon,t}\left[w(t)\left(v_{\phi}(x_{t},t)-\left(\epsilon-x\right)\right)\left(\underbrace{\frac{\partial v_{\phi}(x_{t},t)}{\partial x_{t}}}_{\text{transformer Jacobian}}\frac{\partial x_{t}}{\partial x}+1\right)\frac{\partial x}{\partial\theta}\right]$ (4)
where $\theta$ is the 3D model parameters, $x=g(\mathbf{\theta,c})$ denotes a rendered image from a camera pose $\mathbf{c}$, $\epsilon$ denotes randomly sampled Gaussian noise. Following the convention of the SDS, we omit the transformer Jacobian term for effective training. Therefore, $\left(\frac{\partial v_{\phi}(x_{t},t)}{\partial x_{t}}\frac{\partial x_{t}}{\partial x}+1\right)$ becomes a constant and can be absorbed by $w(t)$, so we have
$\nabla_{\theta}\mathcal{L}_{\mathrm{VFDS}}(\phi,\mathbf{x}=g(\mathbf{\theta,c}))\stackrel{{\scriptstyle\Delta}}{{=}}\mathbb{E}_{\epsilon,t}\left[w(t)\left(v_{\phi}(x_{t},t)-\left(\epsilon-x\right)\right)\frac{\partial x}{\partial\mathbf{\theta}}\right]$ (5)
Based on Equation 5, we train a 3D model utilizing a pretrained rectified flow model. The diffusion model’s trajectory is curved(Lipman et al., 2022), and the score (Song et al., 2020b; Poole et al., 2022) direction varies with different $t$. We denote $\epsilon_{\phi}(x_{t},t)$ as the score function, where $\phi$ represents the parameters of the denoise network, and $x_{t}$ follows the diffusion forward process. In contrast, the trajectory of the rectified flow model (Liu et al., 2022; 2023) is straight, and the vector field direction, $v_{\phi}(x_{t},t)$, remains the same for different $t$. In our VFDS framework – where $t\sim U[0,1]$ is used in every optimization step – the VFDS optimization direction is more consistent. However, the over-smoothing issue of SDS still exists. Therefore, we further analyze the grounding reasons for the over-smoothing issue from the perspective of ODE trajectories.
Elucidating SDS Over-smoothing Issue with VFDS. Because rectified flow is an ODE (Lipman et al., 2022) model, it has the coupling property. Now, we analyze the term $\left(v_{\phi}(x_{t},t)-\left(\epsilon-x\right)\right)$ in Equation 5, where $x_{t}=t\epsilon+(1-t)x$, $\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ is a random sampled noise and $x$ is the generated image from 3D model. As shown in Figure 4, VFDS randomly samples noise $\epsilon$, which leads to multiple ODE trajectories in the same image $x$.
Figure 4: Illustration for over-smoothing analysis. An image is coupled with multiple randomly sampled noises, causing the 3D model to learn ODE trajectories.
When camera poses have only mild differences, the rendered images appear nearly identical. Different ODE trajectories cause inconsistent update directions, which means that directions of $\left(v_{\phi}(x_{t},t)-\left(\epsilon-x\right)\right)$ are inconsistent. Figure 4 illustrates a toy example. $\epsilon^{1},\epsilon^{2},\epsilon^{3}$ are noise randomly sampled from $\mathcal{N}(\mathbf{0},\mathbf{I})$ independently. According to Equation 1, we can get $x_{t}^{1},x_{t}^{2},x_{t}^{3}$, which looks like a blurred image (a hamburger with noise). Following the above analysis, the same $x$ (rightmost hamburger in Figure 4) together with different $\epsilon^{1},\epsilon^{2},\epsilon^{3}$ form different trajectories. During the VFDS training process, our 3D model takes $x_{t}^{1},x_{t}^{2},x_{t}^{3}$ as input and outputs the estimation of trajectory gradients (as shown in Figure 4). Note that the fitting targets are $\epsilon^{1}-x,\epsilon^{2}-x,\epsilon^{3}-x$ respectively. The model finally learns the multiple trajectories gradient, causing the over-smoothing issue.
3.2 Unique Couple Matching Loss
In Sec. 3.1, we have identified the grounding reason for the over-smoothing issue, It is caused by optimizing multiple trajectories during the VFDS training. On top, we propose a novel Unique Couple Matching (UCM) loss that guides the 3D model to optimize in the same trajectory. Our key idea is to leverage the coupling and reversible properties of rectified flow model.
#$\#_{\phi}[x]$ $=x+v_{\phi}(x,\delta_{T_{0}})\delta_{T_{1}}+v_{\phi}(x_{\delta_{T_{1}}},\delta_{T_{1}})\delta_{T_{2}}+\cdots+v_{\phi}(x_{\delta_{T_{n-1}}},\delta_{T_{n-1}})\delta_{T_{n}}$ (6)
$x_{\delta_{T_{n-1}}}$ $=x_{\delta_{T_{n-1}}}+v_{\phi}(x_{\delta_{T_{n-2}}},\delta_{T_{n-2}})\delta_{T_{n-1}},n\geq 2$
where, $\sum_{i=0}^{n}\delta_{T_{i}}=1$, $\delta_{T_{0}}=0$. We define the process from $x$ to $\epsilon$ as push-backward process, denoted ##\##. And #$\#_{\phi}[x]$ denotes iteratively calculate $v_{\phi}(x_{\delta_{T_{i}}},t)$ to backtrack to the $\epsilon$ from $x$ in  Equation 6.
Due to the reversible property of rectified flow, we can search for a noise $\epsilon$ from $x$ in the VFDS framework (see Figure 3(b)). Additionally, because of the aforementioned coupling property, the search noise is unique.
By replacing randomly sampled noise $\epsilon$ to #$\#_{\phi}[x]$ in Equation 5, our UCM loss is defined as follows:
$\nabla_{\theta}\mathcal{L}_{\mathrm{UCM}}(\theta,\mathbf{x}=g(\mathbf{\theta,c}))\stackrel{{\scriptstyle\Delta}}{{=}}\mathbb{E}_{t}\left[w(t)\left(v_{\phi}(x_{t},t)-\left(\#_{\phi}[x]-x\right)\right)\frac{\partial x}{\partial\mathbf{\theta}}\right]$ (7)
$where,x_{t}=t\#_{\phi}[x]+(1-t)x$ (8)
As shown in Figure 3 we can use push-backward instead of randomly sampled noise in VFDS, which guides the 3D model to optimize in the same trajectory.
3.3 FlowDreamer for NeRF and 3D GS
Based on the UCM loss, we propose a novel framework, named FlowDreamer. Our framework yields high-fidelity results with richer textual details and faster convergence. FlowDreamer can be applied to two types of 3D models: 3D Gaussian splatting (Kerbl et al., 2023) and NeRF (Mildenhall et al., 2021).
Application to 3D Gaussian splatting. For 3D Gaussian splatting, we use points generated by the text-to-3D generator (Nichol et al., 2022) as parameter initialization. Then, we directly train the 3D Gaussian splatting model using our UCM loss. Our FlowDreamer yields high-fidelity results with richer textual details (see Figure 1). We also notice our method converges faster than VFDS (see Figure 9).
Application to NeRF and the initialization issue. For NeRF, however, a direct application does not perform well. We call this the initialization issue of NeRF, as we will explain the reasons below. When searching the noise $\epsilon$ for a given image $x$, i.e., push-backward process, we use the rectified flow model $v_{\phi}$. It defines a mapping from data distribution $\pi_{0}$ to noise distribution $\pi_{1}$, where $\pi_{0}$ is the distribution of its pre-training datasets. The effectiveness of the push-backward process depends on that the input distribution for $v_{\phi}$ should be aligned with or at least approximate to $\pi_{0}$. Otherwise, the input lies in an undefined area for $v_{\phi}$ hence the output is unreasonable.
When training NeRF from scratch, it can hardly generate reasonable images based on its randomly initialized parameters. Therefore, the input distribution (denoted as $\pi^{nf}$) is far from $\pi_{0}$, causing the rectified flow model $v_{\phi}$ difficult to estimate the gradient of the ODE trajectory. To solve this issue, we temporarily use the naive VFDS training to warm up as a remedy. We view this initialization issue as an open question and advocate further investigations.
(a) NeRF
(b) 3D Gaussian splatting
Figure 5: Comparison of initialization between NeRF and 3D GS models. (a) Images generated by NeRF models before and after training are different, leading to different push-backward results. (b) As for 3D Gaussian splatting model, the results are quite similar.
As for 3D GS models, the issue does not exist. Taking Figure 5 as an example, when the prompt “A English cottage with stone walls” is provided, the NeRF Initialization is simply a gray image, while the output of the 3D GS model (initialized by Point-E (Nichol et al., 2022)) resembles a cottage. This indicates the initial distribution of the 3D Gaussian splatting model is more approximate to $\pi_{0}$. The result of push-backward process is also more effective, as they are closer to the gradient from images output by the trained model.
4 Experiments
Figure 6: Qualitative comparison under 3D GS generation setting. Our FlowDreamer generates objects with finer details.
4.1 Implementation Details
We adopt Stable Diffusion 3 (SDv3) (Esser et al., 2024) as our Rectified flow model. To facilitate a better comparison, we will categorize the methods into two types: one where the 3D model is NeRF, and the other where the 3D model is 3D GS. For NeRF results comparisons, unlike ProlificDreamer(Wang et al., 2024a) and Consistent3D(Wu et al., 2024), which use multi-stage rendering with normal, geometric, and texture rendering, our method simply uses Instant-NGP (Müller et al., 2022) for rendering. It is optimized with a resolution of 256 for the first 5000 steps with a batch size of 1, and then 512 for the subsequent 3000 steps, also with a batch size of 1. For the warm-up strategy, we use SDS to optimize over 1200 steps. For 3D Gaussian splatting comparisons, we utilize the pretrained PointE (Nichol et al., 2022) to initialize the locations of 3D Gaussians, while other properties of 3D Gaussians adopt random initialization. The batch size and number of iterations are set to 4 and 3000, respectively.
4.2 3D Generation Comparison
Figure 7: Qualitative comparison under NeRF generation setting. Our FlowDreamer generates higher fidelity and sharper results.
NeRF Generation. We compare our FlowDreamer with DreamFusion (Poole et al., 2022), ProlificDreamer (Wang et al., 2024a), Consistent3D (Wu et al., 2024) in NeRF. Other SoTA baselines(Poole et al., 2022; Wang et al., 2024a; Wu et al., 2024) reimplemented by Three-studio(Guo et al., 2023) codebase and employ Stable Diffusion 2.1 for the prior.
As shown in Figure 7, our FlowDreamer achieves results with high fidelity and accurate text alignment. For example, the ‘saguaro cactus’ and ‘clay pot’ exhibit more detail and greater visual quality for the first prompt. Only FlowDreamer does not render the ‘octopus’ and ‘harp’ as a single object for the second prompt.
3D Gaussian splatting Generation. We compare our FlowDreamer with Dreamgaussian(Tang et al., 2023), Gaussiandreamer(Yi et al., 2023) and Luciddreamer(Liang et al., 2023). These 3D GS SoTA baselines are based on their official code by employing the Stable Diffusion 2.1 as the prior.
As shown in Figure 6, our method generates objects that match well with the input text prompts and exhibit realistic textures. For example, our generated ‘pumpkin’ is of high fidelity, and only our method generates the ‘spiders’, which match the text prompt ’plastic’ for the first prompt. The ‘origami pig’ has very rich details, such as its eyes and creases, which are relatively realistic for the second prompt. Our FlowDreamer shows an overall improvement in terms of visual quality and textural details.
4.3 Quantitative Comparisons
Table 1: Quantitative comparisons on CLIP (Radford et al., 2021) similarity with other methods in NeRF generation.
Methods ViT-B-32 ViT-L-14 ViT-g-14
Dreamfusion (Poole et al., 2022) 30.13 29.70 29.49
Prolificdreamer (Wang et al., 2024a) 32.62 32.55 31.60
Consistent3D (Wu et al., 2024) 32.34 32.56 32.01
Ours 34.96 34.19 34.58
We use CLIP (Radford et al., 2021) similarity to quantitatively evaluate our method under either NeRF (Mildenhall et al., 2021) or 3D GS (Kerbl et al., 2023) settings. The results of 3D representation with NeRF come from implementation in  (Guo et al., 2023). The results of 3D representation with 3D GS are from their the official implementation. The prompts of NeRF results are all from Dreamfusion, and the prompts of 3D GS are from LucidDreamer (Liang et al., 2023) and ChatGPT.
Table 2: Quantitative comparisons on CLIP (Radford et al., 2021) similarity with other methods in 3D Gaussian splatting generation.
Methods ViT-B-32 ViT-L-14 ViT-g-14
DreamGaussian (Tang et al., 2023) 22.94 23.50 20.76
GaussianDreamer (Yi et al., 2023) 28.55 29.03 26.98
LucidDreamer (Liang et al., 2023) 28.81 29.78 28.97
Ours 30.70 30.49 30.66
We randomly choose 26 prompts each to compare in 3D GS and NeRF. We randomly select 12 images from the rendered images. The rendered images are from azimuth angles from -180 to 180 degrees with a fixed elevation of 15 degrees for both NeRF and 3D GS. We use three CLIP models from OpenCLIP (Ilharco et al., 2021), ViT-B-32, ViT-L-14, and ViT-g-14, to calculate the CLIP similarity. Our method demonstrates better CLIP similarities both in NeRF and in 3D GS scenarios. As shown in Tab. 1, our CLIP similarity achieved the best results across all three CLIP models, with the largest margin of 2.57 over the second-best result in ViT-g-14 in NeRf results. And Tab. 2 shows that our CLIP similarity also achieved the best results compared to other methods. In particular, it exceeds the LucidDreamer result by 1.89 in ViT-B-32.
4.4 Experimental Insights of our FlowDreamer
Figure 8: Compared to other baseline methods with the same rectified flow prior, our FlowDreamer generates 3D results with richer details and more realistic colors.
3D generation with rectified flow prior. To better demonstrate the effectiveness of our method, we replace the best SOTA method Luciddreamer’s diffusion prior with the Rectified flow prior. The derivation process is provided in the supplementary material. We refer to the ISM loss of LucidDreamer with Rectified Flow Prior as VF-ISM. Figure 8 shows that our method can generate results with finer details and more realistic textures compared with VfDS and VF-ISM. For example, the tank’s tracks and the duck’s feathers are heavily blurred in VFDS. The tank’s color is less realistic, and the color of the duck’s back is somewhat oversaturated in VF-ISM. However, our method not only generates the tank and the duck with richer details but also presents a more realistic appearance.
Figure 9: A comparison of our VFDS and FlowDreamer with different steps. Our method generates images with higher quality than VFDS at the same training steps, indicating a faster convergence speed. Prompt: “A DSLR photo of a large-capacity handbag suitable for travel.”
Faster convergence phenomenon. Our FlowDreamer can be trained to converge faster than our initial framework VFDS. When reaching 1200 steps, our FlowDreamer can generate fine-grained objects whereas VFDS requires more steps to achieve convergence (see Figure 9). When the training step reaches 3000, VFDS still shows much less detail than the FlowDreamer.
Figure 10: A comparison between our VFDS (upper) and FlowDreamer (bottom) with different scales of CFG. Our method shows better generation performance. Prompt: “A cake filled with Oreos, highly detailed, photorealistic.”
Impacts of different Classifier-Free Guidance (CFG) scales. We check the impact of CFG (see Figure 10). The results indicate that we achieve good performance across various CFG scales, demonstrating strong robustness to different CFG scales compared with VFDS.
Figure 11: Impacts of different NFEs. “NFE N+1” denotes using N steps of iteration for push-backward and 1 step for gradient calculation.
Figure 12: Different sampling methods, where ‘NFE: 2*3+1’ indicates that the second-order method requires 2 inferences per iteration, the total push-backward process takes 3 iterations, and an additional evaluation is needed for gradient calculation.
Impacts of different Number of Function Evaluations (NFE). As NFE increases, our FlowDreamer generates objects with more details and more complex structures (see Figure 11). For example, the structure complexity of the front hood of the LEGO car gradually increases. However, Even with a small NFE, wherein the push-backward process has a small cost, our method can still train a 3D model with good performance.
Impacts of various sample methods. We test three sample methods, namely first-order Eulder, second-order midpoint, and fourth-order Runge-Kutta. Experimental results in Figure 12 show that higher-order solvers do not necessarily yield better performance. For more experimental results, please refer to the supplementary materials. For example, the total push-backward process takes three iterations, and the Euler method actually produces more realistic results.
5 Conclusion
In this paper, we explored a new direction by using the rectified flow model as an alternative prior for text-to-3D generation. We developed a mathematical analysis to adapt SDS to rectified flow model, resulting in the initial VFDS framework. However, VFDS still leads to over-smoothing. We analyzed this issue from the perspective of ODE trajectories and proposed FlowDreamer, a text-to-3D framework with a new UCM loss. Extensive experiments showed that FlowDreamer achieves high-fidelity results with richer details and faster convergence in both NeRF and 3D GS settings. We also highlighted open questions, such as initialization issues for NeRF and noise search sampling.
6 Limitation
The Jabus problem still exists; simply adding words like ‘front view,’ ‘back view,’ and ‘side view’ to the prompt is insufficient for supervising the generation view. Although we attempt to mitigate this issue by using Perpneg (Armandpour et al., 2023) in our FlowDreamer method, the problem still occurs occasionally. We consider solving the Jabus problem thoroughly as a focus for future work.